# Preserving Membership Privacy in Low-Rank Adaptation for Latent Diffusion Models

### v2

@@ 4 Experiments
Table 1: Performance of LoRA, MP-LoRA, and SMP-LoRA across all three datasets, as measured by FID, KID, ASR, AUC, and TPR at $5\%$ FPR. Results are presented as mean $\pm$ standard error, based on three independent runs with different seeds. Due to the limited size of the CelebA__\__Small dataset, the FID scores are not available. It is important to note that an AUC value closer to 0.5 (random level) indicates a stronger defensive capability against MI attacks.
@@
4.1 Effectiveness of SMP-LoRA in Defending against MI Attacks
@@
4.2 Ablation Study
This subsection presents ablation studies on the important hyperparameters: the coefficient $\lambda$, the learning rate $\eta_{2}$, and the LoRA’s rank $r$. We also extended the application of SMP-LoRA to the full fine-tuning and DreamBooth (Ruiz et al., 2023) method to assess the generalizability of our membership-privacy-preserving method. Additionally, we compared the performance of SMP-LoRA with gradient clipping and normalization techniques. Furthermore, we evaluated the effectiveness of SMP-LoRA in preserving membership privacy under the black-box MI attacks (Wu et al., 2022) and the white-box gradient-based MI attacks (Pang et al., 2023). Due to space constraints, only the key conclusions are presented in the main paper, with all tables and detailed analyses located in Appendix 6.6.
@@ 4.2 Ablation Study
Table 2 and 6 (Appendix 6.6) present the performance of SMP-LoRA with different coefficient $\lambda\in\{1.00,0.50,0.10,0.05,0.01\}$ across the Pokemon, CelebA__\__Small, and CelebA__\__Large datasets. As $\lambda$ decreases from 1.00 to 0.01, the FID and KID scores gradually decrease, while ASR increases and AUC deviates further from 0.5, suggesting that a lower $\lambda$ shifts the focus more towards minimizing adaptation loss rather than protecting membership privacy.
@@
Figure 3 in Appendix 6.6 shows the ROC curves for SMP-LoRA with these $\lambda$ values across all three datasets. SMP-LoRA effectively defends against MI attacks under both strict False Positive Rate (FPR) constraints and more lenient error tolerance conditions.
Table 2: The effect of coefficient $\lambda$ for SMP-LoRA on the Pokemon dataset. The results for the CelebA__\__Small and CelebA__\__Large datasets can be found in Table 6 (Appendix 6.6).

### v3

@@ A Vanilla Solution: MP-LoRA
Table 1: Performance of LoRA, MP-LoRA, and SMP-LoRA across five datasets, as measured by FID, KID, ASR, AUC, and TPR at $5\%$ FPR. Results are presented as mean $\pm$ standard error, based on three independent runs with different seeds. Due to the limited size of the CelebA__\__Small, CelebA__\__Gender, and CelebA__\__Varying datasets, the FID scores are not available. It is important to note that an AUC value closer to 0.5 (random level) indicates a stronger defensive capability against MI attacks.
@@
Notably, SMP-LoRA also exhibits lower mean and standard deviation of the gradient norm compared to MP-LoRA. We provide further empirical analysis in Appendix D, showing that SMP-LoRA can implicitly rescale the gradient during adaptation by introducing the factors $\frac{1}{1-\lambda G+\delta}$ and $\frac{\mathcal{L}_{\mathrm{ada}}}{\left(1-\lambda G+\delta\right)^{2}}$, thereby leading to more stable gradient and controlled gradient scale compared to MP-LoRA.
4 Experiments
In this section, we first evaluate the performance of LoRA, MP-LoRA, and SMP-LoRA in terms of image generation capability and effectiveness in defending against MI attacks. Then, we conduct ablation studies on the important hyperparameters and further extend our membership-privacy-preserving method to full fine-tuning and DreamBooth (Ruiz et al. 2023) methods. Subsequently, we compare SMP-LoRA with traditional techniques for stabilizing the gradient and evaluate the effectiveness of SMP-LoRA in defending against MI attacks in different settings.
Table 2: Performance of LoRA and SMP-LoRA across three larger datasets, as measured by FID, KID, AUC, and TPR.
TPR
@$5\%$FPR (%) $\downarrow$
CelebA__\__Large__\__5X LoRA 0.53 0.051 92.80 0.94 85.80
SMP-LoRA 0.59 0.062 51.85 0.15 1.70
AFHQ__\__Large__\__5X LoRA 0.39 0.025 88.20 0.86 85.00
SMP-LoRA 0.51 0.041 56.00 0.26 12.00
MS-COCO__\__Large__\__5X LoRA 0.37 0.014 80.40 0.78 68.50
SMP-LoRA 0.72 0.023 46.10 0.14 4.1
Dataset.
In our experiment, we utilized four datasets: Pokemon (Pinkney 2022), CelebA (Liu et al. 2015), AFHQ (Choi et al. 2020), and MS-COCO (Lin et al. 2014). We created several subsets from CelebA, including CelebA__\__Small and CelebA__\__Large, both balanced with equal image contribution per individual, as well as CelebA__\__Gender and CelebA__\__Varying, which are imbalanced with a $7:3$ gender ratio and varied image contributions per individual, respectively. Additiaonlly, we constructed CelebA__\__Large__\__5X, which is five times larger than the CelebA__\__Large along with comparably sized AFHQ__\__Large__\__5X and MS-COCO__\__Large__\__5X. For more details, please refer to Appendix E.
Model hyperparameters.
We utilized the official pre-trained Stable Diffusion v1.5 (CompVis 2022) for LDMs to build the LoRA module, with specific model hyperparameters detailed in Table 5 in Appendix E. We employed a 3-layer MLP as the proxy attack model $h_{\omega}$ and a structurally similar new attack model $h^{\prime}$ to evaluate the effectiveness of adapted LDMs in defending against MI attacks.
Evaluation metrics.
We employed Attack Success Rate (ASR) (Choquette-Choo et al. 2021), Area Under the ROC Curve (AUC), and True Positive Rate (TPR) to evaluate the effectiveness of MP-LoRA and SMP-LoRA in defending against MI attacks. Lower values for ASR and TPR indicate a more effective defense against MI attacks. Consistent with many prior studies (Chen, Yu, and Fritz 2022; Ye et al. 2022; Duan et al. 2023; Dubiński et al. 2024), we default that the TPR measures the capability of the attack model to identify samples as members correctly. Consequently, an AUC value closer to 0.5, i.e., a smaller value of $\frac{\left|\mathrm{AUC}-0.5\right|}{0.5}$, represents a stronger defense capability, as MI attacks involve determining both membership and non-membership. For assessing the image generation capability of the adapted LDMs, we utilized the Fréchet Inception Distance (FID) (Heusel et al. 2017) and the Kernel Inception Distance (KID) (Bińkowski et al. 2018), with lower values denoting better image quality. Please refer to Appendix E for detailed explanations.
Effectiveness of SMP-LoRA in Defending Against MI Attacks
@@ 4 Experiments
In Table 2, we present the performance of LoRA and SMP-LoRA on the CelebA__\__Large__\__5X, AFHQ__\__Large__\__5X, and MS-COCO__\__Large__\__5X datasets. SMP-LoRA remains effective on larger datasets, consistently defending against MI attacks and generating high-quality images.
Ablation Study
This subsection presents ablation studies on the important hyperparameters: the coefficient $\lambda$, the learning rate $\eta_{2}$, and the LoRA’s rank $r$. We also extended the application of SMP-LoRA to the full fine-tuning and DreamBooth (Ruiz et al. 2023) method to assess the generalizability of our membership-privacy-preserving method. Additionally, we compared the performance of SMP-LoRA with gradient clipping and normalization techniques. Furthermore, we evaluated the effectiveness of SMP-LoRA in preserving membership privacy under the black-box MI attacks (Wu et al. 2022) and the white-box gradient-based MI attacks (Pang et al. 2023). Due to space constraints, only the key conclusions are presented in the main paper, with all tables and detailed analyses located in Appendix F.
@@
Table 6 (Appendix F) presents the performance of SMP-LoRA with different coefficient $\lambda\in\{1.00,0.50,0.10,0.05,0.01\}$ across the Pokemon, CelebA__\__Small, and CelebA__\__Large datasets. As $\lambda$ decreases from 1.00 to 0.01, the FID and KID scores gradually decrease, while ASR increases and AUC deviates further from 0.5, suggesting that a lower $\lambda$ shifts the focus more towards minimizing adaptation loss rather than protecting membership privacy.
@@
Figure 3 in Appendix F shows the ROC curves for SMP-LoRA with these $\lambda$ values across all three datasets. SMP-LoRA effectively defends against MI attacks under both strict False Positive Rate (FPR) constraints and more lenient error tolerance conditions.
Learning rate $\eta_{2}$.
Table 7 in Appendix F displays the performance of SMP-LoRA with different learning rates $\eta_{2}\in\{1e-4,1e-5,1e-6\}$ on the Pokemon dataset. SMP-LoRA consistently preserves membership privacy across all tested learning rates.
LoRA’s rank $r$.
Table 8 in Appendix F shows the performance of SMP-LoRA with different rank $r\in\{128,64,32,16,8\}$ on the Pokemon dataset. The performance of SMP-LoRA is not significantly affected by the LoRA’s rank $r$.
Extending to the full fine-tuning and DreamBooth Methods.
In Table 9 (Appendix F), we present the performance of SMP-LoRA and its extension to the full fine-tuning and DreamBooth (Ruiz et al. 2023) methods on the Pokemon dataset. Our membership-privacy-preserving method continues to effectively protect membership privacy when applied to these methods, highlighting its potential applicability across different adaptation methods.
Comparing with gradient clipping and normalization techniques.
In Table 10 (Appendix F), we report the performance of SMP-LoRA and MP-LoRA enhanced with gradient clipping and normalization techniques on the Pokemon dataset. These traditional techniques for stabilizing the gradient cannot address the unstable optimization issue in MP-LoRA.
Defending against MI attacks in different settings.
In Table 11 (Appendix F), we display the attack performance on LoRA and SMP-LoRA using the black-box MI attack (Wu et al. 2022) and the white-box gradient-based MI attack (Pang et al. 2023), which is currently the most potent MI attack targeting DMs. This renders further comparisons with weaker attacks unnecessary, such as gray-box MI attacks (Duan et al. 2023; Kong et al. 2024; Fu et al. 2023). Compared to LoRA, SMP-LoRA, specifically designed to defend against white-box loss-based MI attacks, consistently provides enhanced membership privacy protection when facing MI attacks in different settings. Implementation details for the black-box and white-box gradient-based MI attacks are available in Appendix G.
5 Conclusion
Acknowledgments
Feng Liu is supported by the Australian Research Council (ARC) with grant numbers DP230101540 and DE240101089, and the NSF&CSIRO Responsible AI program with grant number 2303037.
Di Wang is supported in part by the funding BAS/1/1689-01-01, URF/1/4663-01-01, REI/1/5232-01-01, REI/1/5332-01-01, and URF/1/5508-01-01 from KAUST, and funding from KAUST - Center of Excellence for Generative AI, under award number 5940.
Appendix A Algorithm of MP-LoRA
We provide the specific implementation of MP-LoRA as follows:
Algorithm 2 Membership-Privacy-preserving LoRA
Output: a MP-LoRA for LDMs
1:  Perform low-rank decomposition on $f_{\theta}$ to obtain $f_{\bar{\theta}+\mathbf{B}\mathbf{A}}$ ($\mathbf{B}$ and $\mathbf{A}$ are trainable LoRA modules)
2:  for each epoch do
3:     for each training iteration do
4:        Sample batches $S^{\mathrm{m}}$ and $S^{\mathrm{nm}}$ from $\mathcal{D}^{\mathrm{m}}_{\mathrm{aux}}$ and $\mathcal{D}^{\mathrm{nm}}_{\mathrm{aux}}$
5:        Calculate the MI gain $G^{*}$ on $S^{\mathrm{m}}\cup S^{\mathrm{nm}}$
6:        Update the parameters $\omega\leftarrow\omega+\eta_{1}\cdot\nabla_{\omega}G^{*}$.
7:        Sample a fresh batch from $\mathcal{D}_{\mathrm{tr}}$
8:        Calculate the training loss $\mathcal{L}^{*}=\mathcal{L}_{\mathrm{PL}}$
9:        Update parameters $\mathbf{A}\leftarrow\mathbf{A}-\eta_{2}\cdot\nabla_{\mathbf{A}}\mathcal{L}^{*}$ and $\mathbf{B}\leftarrow\mathbf{B}-\eta_{2}\cdot\nabla_{\mathbf{B}}\mathcal{L}^{*}$, respectively
10:     end for
11:  end for
(a) Gradient scale obtained by training loss.
(b) Gradient scale of MP-LoRA.
(c) Gradient scale of SMP-LoRA.
Figure 2: Figure 2(a) shows the mean and standard deviation of gradient scales obtained by training loss throughout the training iterations of LoRA, MP-LoRA, and SMP-LoRA on the Pokemon dataset. Figure 2(b) and 2(c) display the gradient scales obtained by the adaptation loss and MI gain respectively during MP-LoRA and SMP-LoRA. Compared with MP-LoRA, SMP-LoRA has more stable gradient and controlled gradient scale.
Appendix B Detailed Derivation for Establishing the Relaxed Smoothness Condition
We provide the detailed derivation for establishing the relaxed smoothness condition for MP-LoRA and SMP-LoRA, respectively.
For MP-LoRA, we derive the second derivatives of the training loss $\mathcal{L}_{\mathrm{PL}}$ in Equation (6) to establish the relaxed smoothness condition.
$\frac{\partial^{2}\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}^{2}}$ $=\frac{\partial}{\partial\mathbf{B}\mathbf{A}}\left(\frac{\partial\mathcal{L}_{\mathrm{PL}}}{\partial\mathbf{B}\mathbf{A}}\right)$